# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v2

@@ 6.3. The 2-Sphere: Trivial Fundamental Group
Each letter in $w$ comes from either $\pi_{1}(A)$ or $\pi_{1}(B)$, both trivial. Thus each letter represents the identity loop $\rho]$, and $\mathrm{decode}(w)=[\rho]^{n}=[\rho]$. ∎
@@
1theorem trivial_decode (f : C -> PUnit’) (g : C -> PUnit’) (c0 : C)
2 (w : FreeProductWord (pi_1(PUnit’, PUnit’.unit)) (pi_1(PUnit’, PUnit’.unit))) :
3 pushoutDecode c0 w = Quot.mk _ (Path.refl _) := by
4 induction w with
5 | nil => rfl
6 | consLeft alpha rest ih =>
7 simp only [pushoutDecode]
8 rw [trivial_left_inclusion c0 alpha]
9 rw [ih]
10 exact piOneMul_refl_left _
11 | consRight beta rest ih =>
12 simp only [pushoutDecode]
13 rw [trivial_right_inclusion c0 beta]
14 rw [ih]
15 exact piOneMul_refl_left _
16
17theorem sphere2_pi1_trivial :
18 forall (alpha : pi_1(Sphere2, basepoint)), alpha = Quot.mk _ (Path.refl _) := ...
19
20noncomputable def sphere2_pi1_equiv_unit :
21 SimpleEquiv (pi_1(Sphere2, basepoint)) Unit where
22 toFun := fun _ => ()
23 invFun := fun _ => Quot.mk _ (Path.refl _)
24 left_inv := by intro alpha; exact (sphere2_pi1_trivial alpha).symm
25 right_inv := by intro u; cases u; rfl
Remark 6.5 (Higher Homotopy Groups).
The SVK theorem has no direct analogue for higher homotopy groups: $\pi_{n}(A\vee B)\not\simeq\pi_{n}(A)\times\pi_{n}(B)$ in general for $n>1$. Instead, the Blakers-Massey theorem [8] provides connectivity bounds. Computing higher $\pi_{n}$ via computational paths would require the weak $\omega$-groupoid structure established in prior work [5].
6.4. The 3-Sphere and the Hopf Fibration
The SVK technique for proving $\pi_{1}(S^{2})\simeq 1$ generalizes immediately to higher-dimensional spheres. We demonstrate this with the 3-sphere.
Theorem 6.6 (Fundamental Group of 3-Sphere).
$\pi_{1}(S^{3})\simeq 1$
@@ 6.4. The 3-Sphere and the Hopf Fibration
The 3-sphere is the suspension of the 2-sphere: $S^{3}=\Sigma(S^{2})$.
As a pushout: $\Sigma(S^{2})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{2},!,!)$.
By SVK:
$\pi_{1}(S^{3})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{2})}\pi_{1}(\mathbf{1})=1\ast_{1}1=1$
The proof is identical in structure to Theorem 6.3—only the amalgamating group changes from $\pi_{1}(S^{1})\cong\mathbb{Z}$ to $\pi_{1}(S^{2})\cong 1$. Since both factors $\pi_{1}(\mathbf{1})$ are trivial, the result is trivial regardless. ∎
This pattern continues: for any $n\geq 2$, we have $S^{n}=\Sigma(S^{n-1})$, and by induction $\pi_{1}(S^{n})\simeq 1$.
In Lean, the proof reuses the same SVK machinery:
1theorem sphere3_pi1_trivial :
2 forall (l : LoopSpace Sphere3 sphere3North),
3 Quot.mk RwEq l = Quot.mk RwEq (Path.refl sphere3North) := by
4 intro l
5 let f : Sphere2 -> PUnit := sphere2ToNorth
6 let g : Sphere2 -> PUnit := sphere2ToSouth
7 let c0 : Sphere2 := basepoint
8 let encoded := pushoutEncodeAmalg c0 (Quot.mk RwEq l)
9 have left_inv_l : Quot.mk RwEq l = pushoutDecodeAmalg c0 encoded := by
10 simp only [pushoutDecodeAmalg]
11 exact (pushoutDecodeEncodeAxiom PUnit PUnit Sphere2 f g c0 l).symm
12 rw [left_inv_l]
13 exact amalg_trivial_is_one_s3 encoded
6.4.1. The Hopf Fibration Context
The 3-sphere is the total space of the Hopf fibration, a fundamental fiber bundle in topology:
$S^{1}\longrightarrow S^{3}\longrightarrow S^{2}$
This bundle has fiber $S^{1}$, total space $S^{3}$, and base $S^{2}$.
In the computational paths framework, we model this via a type family:
  • Hopf fiber family: $P:S^{2}\to\mathcal{U}$ with $P(x)=S^{1}$ for all $x$
  • Total space: $\Sigma_{x:S^{2}}P(x)$, equivalent to $S^{3}$
  • Hopf projection: $\mathrm{hopfProj}:\Sigma_{x:S^{2}}P(x)\to S^{2}$
The fiber equivalence $\mathrm{Fiber}(\mathrm{hopfProj},x)\simeq S^{1}$ is proved (not axiomatized) from the $\Sigma$-type construction.
Remark 6.7 (Long Exact Sequence).
The Hopf fibration induces a long exact sequence of homotopy groups:
$\cdots\to\pi_{2}(S^{1})\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\pi_{1}(S^{1})\to\pi_{1}(S^{3})\to\pi_{1}(S^{2})\to\cdots$
Substituting known values:
$\cdots\to 1\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\mathbb{Z}\to 1\to 1$
By exactness, the connecting map $\partial:\pi_{2}(S^{2})\to\mathbb{Z}$ is an isomorphism (in untruncated homotopy theory), yielding $\pi_{2}(S^{2})\cong\mathbb{Z}$. In our truncated framework where $\pi_{2}\cong 1$, this collapses, but the fibration structure remains valuable for understanding the topology.
Remark 6.8 (Non-Triviality).
The Hopf fibration is non-trivial: $S^{3}\not\simeq S^{2}\times S^{1}$. If it were a trivial product, we would have $\pi_{1}(S^{3})\cong\pi_{1}(S^{2})\times\pi_{1}(S^{1})\cong 1\times\mathbb{Z}\cong\mathbb{Z}$, contradicting $\pi_{1}(S^{3})\cong 1$.
@@ 7.1. Architecture
Path/Rewrite/Step.lean Single-step rewrite rules (50+)
Path/Rewrite/RwEq.lean Rewrite equality relation
@@
Path/HIT/Pushout.lean Pushout HIT axioms
@@
Path/HIT/FigureEight.lean Figure-eight application
Path/HIT/Sphere.lean 2-sphere application
Path/HIT/Torus.lean $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$
Path/HIT/KleinBottle.lean $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$
Path/HIT/OrientableSurface.lean $\pi_{1}(\Sigma_{g})$ for genus $g$
Path/HIT/HopfFibration.lean $S^{3}$, Hopf map, $\pi_{1}(S^{3})=1$
Path/Homotopy/Fibration.lean Fiber sequences, long exact sequence
Path/Homotopy/EilenbergMacLane.lean $K(G,n)$ spaces
7.2. Key Definitions
The formalization uses Lean 4’s quotient types for both $\pi_{1}$ (quotient by $\sim$) and the amalgamated free product (quotient by amalgamation equivalence).
1-- Fundamental group as quotient
2def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
@@ 7.2. Key Definitions
4-- Amalgamated free product as quotient
5def AmalgamatedFreeProduct (G1 G2 H : Type u)
6 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
7 Quot (AmalgEquiv i1 i2)
7.3. Axiom Inventory
Table 1. Axiom usage in the formalization
Axiom Purpose Justification
Pushout Pushout type exists Standard HIT
Pushout.inl/inr Point constructors Standard HIT
Pushout.glue Path constructor Standard HIT
glue_natural_rweq Naturality of glue Derivable in cubical TT
pushoutEncodeAxiom Encode function exists Code family construction
pushoutEncodeAxiom_respects_rweq Encode respects $\sim$ From encode construction
pushoutDecodeEncodeAxiom Round-trip property Path induction
pushoutEncodeDecodeAxiom Round-trip property Word induction
Remark 7.1 (Noncomputability).
Definitions involving HITs are marked noncomputable because Lean 4 lacks native computational support for higher-inductive types. The decode function, while marked noncomputable, has constructive content—given a word, it produces an explicit loop term. The encode function is genuinely nonconstructive in our axiomatization.
Remark 7.2 (Consistency).
The pushout axioms are standard HIT axioms known to be consistent with MLTT via various models (simplicial sets, cubical sets [1]). The encode axioms are consequences of the full HIT elimination principle, which is derivable in cubical type theories.
7.4. What Is Proved vs. Axiomatized
Table 2. Proof status of main results
Result Status Dependencies
Decode function Proved None
Decode respects amalgamation Proved glue_natural_rweq
Encode function Axiomatized pushoutEncodeAxiom
Round-trip ($\mathrm{decode}\circ\mathrm{encode}$) Axiomatized pushoutDecodeEncodeAxiom
Round-trip ($\mathrm{encode}\circ\mathrm{decode}$) Axiomatized pushoutEncodeDecodeAxiom
SVK equivalence Proved Above axioms
Wedge sum theorem Proved SVK + triviality
Figure-eight $\pi_{1}$ Proved Wedge + $\pi_{1}(S^{1})\simeq\mathbb{Z}$
2-sphere $\pi_{1}$ Proved SVK + triviality
3-sphere $\pi_{1}$ Proved SVK + triviality (same as $S^{2}$)
Non-abelianness of $\mathbb{Z}*\mathbb{Z}$ Proved Definitional (constructor ineq.)

### v3

@@ 6.6. The Real Projective Plane
$\mathbb{RP}^{2}$ can be constructed by attaching a disk to a circle via the map $z\mapsto z^{2}$. The fundamental group has one generator (the loop) with the relation that going around twice is contractible. ∎
6.7. Orientable Surfaces of Genus $g$
Theorem 6.10 (Fundamental Group of Orientable Surface).
$\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid[a_{1},b_{1}]\cdots[a_{g},b_{g}]=1\rangle$
where $a,b]=aba^{-1}b^{-1}$ is the commutator.
@@ 6.7. Orientable Surfaces of Genus $g$
1def OrientableSurfaceWord (g : Nat) : Type :=
2 List (Fin (2 * g) * Bool) -- generator index and sign
3
4inductive SurfaceRelation (g : Nat) :
5 OrientableSurfaceWord g -> OrientableSurfaceWord g -> Prop
6 | commutator_product :
7 -- [a_1, b_1] ... [a_g, b_g] = 1
8 SurfaceRelation g (commutatorProduct g) []
6.8. Non-Orientable Surfaces
Theorem 6.11 (Fundamental Group of Non-Orientable Surface).
$\pi_{1}(N_{n})\simeq\langle a_{1},\ldots,a_{n}\mid a_{1}^{2}\cdots a_{n}^{2}=1\rangle$
Example 6.12 (Special Cases).
  • $N_{1}=\mathbb{RP}^{2}$: $\pi_{1}(N_{1})=\langle a\mid a^{2}=1\rangle\cong\mathbb{Z}_{2}$
  • $N_{2}=K$ (Klein bottle): $\pi_{1}(N_{2})=\langle a,b\mid a^{2}b^{2}=1\rangle\cong\mathbb{Z}\rtimes\mathbb{Z}$
6.9. Lens Spaces
Theorem 6.13 (Fundamental Group of Lens Space).
For coprime $p,q$:
$\pi_{1}(L(p,q))\simeq\mathbb{Z}/p\mathbb{Z}$
Remark 6.14 (Lens Space Construction).
Lens spaces $L(p,q)$ are quotients of $S^{3}$ by cyclic group actions. The fundamental group depends only on $p$, not $q$; however, the spaces $L(p,q)$ and $L(p,q^{\prime})$ may be non-homeomorphic for different values of $q$.
6.10. Bouquet of $n$ Circles
Theorem 6.15 (Fundamental Group of $n$-Bouquet).
$\pi_{1}(\bigvee_{n}S^{1})\simeq F_{n}$
where $F_{n}$ is the free group on $n$ generators.
7. Extended Features
7.1. Higher Homotopy Groups
Definition 7.1 (Higher Homotopy Groups).
The $n$-th homotopy group is defined via iterated loop spaces:
$\pi_{n}(A,a):=\pi_{1}(\Omega^{n-1}(A,a),\rho^{n-1})$
where $\Omega^{n}$ denotes $n$-fold iteration of the loop space.
Theorem 7.2 (Eckmann-Hilton).
For $n\geq 2$, $\pi_{n}(A,a)$ is abelian.
@@ 7.1. Higher Homotopy Groups
The Eckmann-Hilton argument shows that when two binary operations share a common unit and satisfy an interchange law, they must coincide and be commutative. In $\Omega^{2}(A,a)$, both horizontal and vertical composition of 2-loops satisfy these conditions. ∎
In Lean:
1def Loop2Space (A : Type u) (a : A) : Type u :=
2 Derivation_2 (Path.refl a) (Path.refl a)
3
4def PiTwo (A : Type u) (a : A) : Type u :=
5 Quotient (Loop2Setoid A a)
6
7theorem piTwo_comm (x y : PiTwo A a) : PiTwo.mul x y = PiTwo.mul y x := by
8 induction x using Quotient.ind
9 induction y using Quotient.ind
10 apply Quotient.sound
11 exact contractibility_3 _ _
7.2. The Weak $\omega$-Groupoid Structure
Theorem 7.3 (Types are Weak $\omega$-Groupoids).
Via computational paths, every type carries the structure of a weak $\omega$-groupoid:
  • 0-cells: points of AA
  • 1-cells: paths $p:\mathrm{Path}(a,b)$
  • 2-cells: derivations $\mathcal{D}_{2}(p,q)$ witnessing $p\sim q$
  • 3-cells: meta-derivations $\mathcal{D}_{3}(\alpha,\beta)$
  • $n$-cells: $\mathcal{D}_{n}$ for all $n$
In Lean:
1-- 2-cells: derivations between paths
2inductive Derivation_2 : Path a b -> Path a b -> Type u
3 | refl (p : Path a b) : Derivation_2 p p
4 | step (s : Step p q) : Derivation_2 p q
5 | inv (d : Derivation_2 p q) : Derivation_2 q p
6 | vcomp (d1 : Derivation_2 p q) (d2 : Derivation_2 q r) : Derivation_2 p r
7
8-- 3-cells: derivations between 2-cells
9inductive Derivation_3 : Derivation_2 p q -> Derivation_2 p q -> Type u
10 | refl (d : Derivation_2 p q) : Derivation_3 d d
11 | step (m : MetaStep_2 d1 d2) : Derivation_3 d1 d2
12 | inv (m : Derivation_3 d1 d2) : Derivation_3 d2 d1
13 | vcomp (m1 : Derivation_3 d1 d2) (m2 : Derivation_3 d2 d3) : Derivation_3 d1 d3
7.3. Truncation Levels
Definition 7.4 (Truncation Levels).
Following HoTT, we define truncation levels:
  • IsContr(AA): AA is contractible (has a unique point)
  • IsProp(AA): any two points are equal (“propositions”)
  • IsSet(AA): any two parallel paths are equal (“sets”)
  • IsGroupoid(AA): any two parallel 2-cells are equal
In Lean:
1class IsContr (A : Type u) where
2 center : A
3 contr : forall a, Path center a
4
5class IsProp (A : Type u) where
6 allEq : forall (a b : A), Path a b
7
8class IsSet (A : Type u) where
9 pathEq : forall (a b : A) (p q : Path a b), RwEq p q
10
11class IsGroupoid (A : Type u) where
12 derivEq : forall (a b : A) (p q : Path a b) (d e : Derivation_2 p q),
13 Derivation_3 d e
7.4. Covering Space Theory
Definition 7.5 (Covering Space).
A covering space of $B$ with fiber $F$ consists of:
  • A total space $E$
  • A projection $p:E\to B$
  • A $\pi_{1}(B)$-action on $F$
  • Fiber equivalence: $p^{-1}(b)\simeq F$ for each $b$
In Lean:
1structure CoveringSpace (B : Type u) (F : Type u) (b0 : B) where
2 totalSpace : Type u
3 proj : totalSpace -> B
4 fiber_equiv : forall b, Equiv (Fiber proj b) F
5 deck_action : pi_1(B, b0) -> Equiv F F
6 action_coherent : -- action is compatible with path lifting
7.5. Fibration Theory
Definition 7.6 (Fibration).
A fibration $p:E\to B$ with fiber $F$ admits a long exact sequence:
$\cdots\to\pi_{2}(F)\to\pi_{2}(E)\to\pi_{2}(B)\to\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)\to 1$
7.6. Eilenberg-MacLane Spaces
Theorem 7.7 ($S^{1}$ as $K(\mathbb{Z},1)$).
The circle is the Eilenberg-MacLane space $K(\mathbb{Z},1)$:
  • $\pi_{1}(S^{1})\cong\mathbb{Z}$
  • $\pi_{n}(S^{1})\cong 1$ for $n\geq 2$
7.7. Path Simplification Tactics
The formalization includes automated tactics for path reasoning:
1-- Basic simplification using groupoid laws
2path_simp -- refl . p ~ p, p . refl ~ p, etc.
3
4-- Full automation (~25 simp lemmas)
5path_auto -- handles most standalone RwEq goals
6
7-- Specialized tactics
8path_rfl -- close p ~ p
9path_canon -- normalize to canonical form
10path_cancel_left -- p^-1 . p ~ refl
11path_cancel_right -- p . p^-1 ~ refl
12path_congr_left h -- apply hypothesis on right
13path_congr_right h -- apply hypothesis on left
@@ 8. The Lean 4 Formalization
8.1. Why Axioms Are Necessary
Lean 4 does not natively support higher-inductive types (HITs). Unlike Agda’s --cubical mode or Coq’s HoTT library, there is no built-in mechanism for defining types with both point and path constructors. We therefore axiomatize HITs, introducing axioms only where strictly necessary.
Definition 8.1 (Kernel Axiom).
A kernel axiom is a Lean axiom declaration that extends the trusted kernel. These are used sparingly for:
  1. (1)
    HIT type declarations: The type itself (e.g., Circle : Type u)
  2. (2)
    Point constructors: Base points (e.g., circleBase : Circle)
  3. (3)
    Path constructors: Generating paths (e.g., circleLoop : Path circleBase circleBase)
  4. (4)
    Higher path constructors: 2-cells, surface relations
  5. (5)
    Recursion principles: Eliminators with $\beta$-rules
Definition 8.2 (Typeclass Assumption).
A typeclass assumption is a hypothesis provided via Lean’s typeclass mechanism. These are not kernel axioms—they are parameters that users instantiate. We use these for:
  1. (1)
    Computation rules: HasRecGlueRwEq (how rec acts on glue)
  2. (2)
    Naturality: HasGlueNaturalLoopRwEq (glue commutes with paths)
  3. (3)
    Encode/decode data: HasPushoutSVKEncodeQuot, etc.
Remark 8.3 (Axioms vs Assumptions).
The distinction is crucial:
  • Kernel axioms (36 total) are permanently trusted by Lean. We minimize these.
  • Typeclass assumptions are user-provided hypotheses. They appear in theorem statements as [HasFoo] parameters, making dependencies explicit.
For example, seifertVanKampenEquiv requires several typeclass assumptions but introduces no new kernel axioms—users must provide evidence for the assumptions.
Remark 8.4 (Consistency of the Encode/Decode Axioms).
A natural question arises: why doesn’t the fundamental group collapse? In Lean’s proof-irrelevant Prop, all equality proofs between the same terms are equal. If $\pi_{1}$ depended only on equality proofs, all loops would be identified.
The key to consistency is threefold:
  1. (1)
    Explicit step lists (see Remark 2.1): Paths store explicit steps : List (Step A), and $\sim$ identifies paths only when their step sequences can be transformed via the Step rules—not merely when their underlying equality proofs coincide.
  2. (2)
    No path collapse rule: The Step relation contains no rule that identifies arbitrary paths with the same endpoints. In particular, there is no “canonicalization” rule $p\rightsquigarrow\mathrm{ofEq}(p.\mathrm{proof})$ that would reduce any path to a canonical form based solely on its equality proof.
  3. (3)
    Axiomatized classification data: The encode/decode equivalences (e.g., $\pi_{1}(S^{1})\simeq\mathbb{Z}$) are assumed via typeclass assumptions, not derived. We do not assert that transport along a self-equality computes nontrivially—such an assertion would be inconsistent with Lean’s proof-irrelevant Prop. Instead, we assume the classification result: there exists an equivalence between $\pi_{1}$ and the expected group, with the round-trip properties holding.
This design follows the HoTT tradition of axiomatizing what cannot be computed in the ambient type theory. The typeclass assumptions for encode/decode are analogous to Licata-Shulman’s encode-decode method [10], but formulated as external assumptions rather than internal proofs.
8.2. Justifying the 36 Kernel Axioms
Each HIT requires axioms because Lean cannot construct them from primitive types:
HIT Why Axioms Are Needed
Circle Needs a non-trivial loop $\mathrm{loop}:\mathrm{Path}(\mathrm{base},\mathrm{base})$ that is not $\rho$. No Lean type has this property without axioms.
Sphere Needs 2-dimensional structure (meridians collapsing at poles).
Klein bottle Needs the non-orientable surface relation $aba^{-1}b=1$.
$\mathbb{RP}^{2}$ Needs $a^{2}=1$ as a path relation.
$\Sigma_{g}$, $N_{n}$ Need genus-dependent relations on generators.
Cylinder Needs both a loop and a segment with a 2-cell between them.
Remark 8.5 (Axiom-Free Constructions).
Some spaces require no additional kernel axioms beyond those already listed:
  • Torus: Defined as $T^{2}:=S^{1}\times S^{1}$ (uses Circle axioms only)
  • Möbius band: Homotopy equivalent to $S^{1}$ (uses Circle axioms only)
  • Wedge sum: Defined as $\mathrm{Pushout}(A,B,\mathbf{1})$ (zero kernel axioms—Pushout is Quot-based)
  • Suspension: Defined as $\mathrm{Pushout}(\mathbf{1},\mathbf{1},A)$ (zero kernel axioms—Pushout is Quot-based)
  • Sphere: Defined as $S^{n}:=\Sigma^{n}(\mathbf{1})$ via iterated suspension (zero kernel axioms)
8.3. When to Use Each SVK Assumption
The split typeclass architecture allows users to provide minimal hypotheses:
Assumption When to Use
HasGlueNaturalLoopRwEq Always required for SVK. Automatically derived when $C$ is a subsingleton (wedge) or when $A$, $B$ satisfy Axiom K.
HasPushoutSVKEncodeQuot When you have an explicit encode function.
HasPushoutSVKDecodeEncode Standard encode-decode round-trip.
HasPushoutSVKEncodeDecode Round-trip up to amalgamation equivalence.
HasPushoutSVKDecodeAmalgBijective Choice-based interface when you only know decode is bijective (Prop-level), without explicit encode.
@@ 8.4. Architecture
Path/Basic/Congruence.lean congrArg, transport
Path/Rewrite/Step.lean Single-step rewrites (50+ rules)
Path/Rewrite/RwEq.lean Rewrite equality
Path/Rewrite/PathTactic.lean Automated tactics
Path/Homotopy/Loops.lean Loop spaces
@@
Path/Homotopy/HigherHomotopy.lean $\pi_{n}$ for $n\geq 2$
Path/HIT/Circle.lean Circle HIT
Path/HIT/Pushout.lean Pushout HIT
@@
Path/HIT/FigureEight.lean $S^{1}\vee S^{1}$
Path/HIT/OrientableSurface.lean $\Sigma_{g}$
Path/HIT/NonOrientableSurface.lean $N_{n}$
Path/HIT/KleinBottle.lean Klein bottle
Path/HIT/ProjectivePlane.lean $\mathbb{RP}^{2}$
Path/HIT/LensSpace.lean $L(p,q)$
Path/OmegaGroupoid.lean Weak $\omega$-groupoid structure
8.5. Kernel Axiom Inventory
The formalization uses exactly 36 kernel axioms, all for higher-inductive types:
Table 2. Kernel axioms by HIT
HIT Axioms Description
Circle 7 Type, base, loop, rec, rec_base, rec_loop, ind
[…]
KleinBottle 5 Type, base, loopA, loopB, surface
OrientableSurface 5 Type, base, loopA, loopB, surfaceRelation
NonOrientableSurface 4 Type, base, loop, 2-cell
ProjectivePlane 4 Type, base, loop, loopSquare
Total 36
Notable non-axioms (constructions that add 0 kernel axioms):
  • Pushout: Defined as Quot (PushoutRel A B C f g) using Lean’s built-in quotient type—no kernel axioms. Constructors inl, inr, glue are definitions, not axioms.
  • Sphere: Defined as $S^{n}:=\Sigma^{n}(\mathbf{1})$ via suspension of the unit type—uses Pushout.
  • Torus: Constructed as $S^{1}\times S^{1}$—uses only Circle axioms.
  • Möbius band: Homotopy equivalent to $S^{1}$—uses only Circle axioms.
  • Wedge, Suspension: Defined as special cases of Pushout.
  • SVK encode/decode: Typeclass assumptions, not kernel axioms.
  • Lens spaces, Bouquet: Constructed from existing HITs.
Remark 8.6 (Pushout Design).
Unlike HITs such as Circle or Klein bottle (which require kernel axioms for their path constructors), the Pushout is implemented using Lean’s built-in Quot type. The constructors inl, inr are Quot.mk, and glue is Path.ofEq (Quot.sound ...).
The computation rules (how rec acts on glue, glue naturality) are provided via typeclass assumptions (HasRecGlueRwEq, HasGlueNaturalLoopRwEq), not kernel axioms. This design means Pushout adds zero kernel axioms; all structure comes from user-instantiated typeclasses.
8.6. Opt-in Axiom Imports
For convenience, opt-in imports provide assumption-free APIs:
1-- Import to get circlePiOneEquivInt’ with no hypotheses
2import ComputationalPaths.Path.HIT.CirclePiOneAxiom
@@ 8.6. Opt-in Axiom Imports
4-- Import to get torusPiOneEquivIntProd’ with no hypotheses
5import ComputationalPaths.Path.HIT.TorusPiOneAxiom
6
7-- Import to get wedgeFundamentalGroupEquiv’ with no hypotheses
8import ComputationalPaths.Path.HIT.WedgeSVKAxiom